OpenStudy (anonymous):

Two competing cell phone companies have recently restructured their payment plans. Mobilama requires a $20 activation fee and charges $25 per month for their basic plan. Ringit Wireless requires no activation fee and charges $35 per month for their basic plan. When will the two plans be equal in cost?

OpenStudy (anonymous):

After 2 months Ringit = 35 *2= 70 and after 2 months Mobilama 25 *2 +20= 70 hope this helped your @arishagee but its jst a hit and trial one...

OpenStudy (anonymous):

thank you!!